OpenStudy (anonymous):

Solve. 2c/ c-4 -2 = 4/c+5

OpenStudy (anonymous):

\[\frac{ 2c }{ c-4 } -2 =\frac{ 4 }{ c+5 }\]

OpenStudy (anonymous):

On the left side, 2c/c the c cancels out because the ratio is 1. Therefore, you have (2/-4)-2 on the left side. Next you could simplify the left side to -2-2 and then further to -4. Next, you can multiply both sides by c+5. You will then be left with -4(c+5)=4. Now you distribute the -4 and then isolate the variable.

OpenStudy (anonymous):

I'm confused. How do I distribute the -4 and isolate the variable?

OpenStudy (anonymous):